# MindGYM: Enhancing Vision-Language Models via Synthetic Self-Challenging Questions

### v1

@@ Abstract
Large vision-language models (VLMs) face challenges in achieving robust, transferable reasoning abilities due to reliance on labor-intensive manual instruction datasets or computationally expensive self-supervised methods. To address these issues, we introduce MindGYM, a framework that enhances VLMs through synthetic self-challenging questions, consisting of three stages: (1) Seed Single-Hop Question Synthesis, generating cognitive questions across textual (e.g., logical deduction) and multimodal contexts (e.g., diagram-based queries) spanning eight semantic areas like ethical analysis; (2) Challenging Multi-Hop Question Synthesis, combining seed questions via diverse principles like bridging, visual-textual alignment, to create multi-step problems demanding deeper reasoning; and (3) Thinking-Induced Curriculum Fine-Tuning, a structured pipeline that progressively trains the model from scaffolded reasoning to standalone inference. By leveraging the model’s self-synthesis capability, MindGYM achieves high data efficiency (e.g., +16% gains on MathVision-Mini with only 400 samples), computational efficiency (reducing both training and inference costs), and robust generalization across tasks. Extensive evaluations on seven benchmarks demonstrate superior performance over strong baselines, with notable improvements (+15.77% win rates) in reasoning depth and breadth validated via GPT-based scoring. MindGYM underscores the viability of self-challenging for refining VLM capabilities while minimizing human intervention and resource demands. Code and data are released to advance multimodal reasoning research.
**footnotetext: Equal contributions.
@@ 1 Introduction
Large vision-language models (VLMs) have emerged as foundational tools for tasks requiring joint reasoning over visual and textual inputs [4, 30]. However, enabling robust performance with transferable and efficient reasoning abilities with high-quality data remains challenging [6, 8]. Manually curated instruction datasets like OK-VQA [25] and ScienceQA [22] are labor-intensive to scale, while self-supervised synthetics methods such as MMInstruct [21] and MMEvol [24] suffer from limited generalization cross modality and tasks, failing to tightly and logically couple visual and textual semantics. Meanwhile, reasoning enhancement methods—such as reinforcement learning (RL) [1] or iterative prompting [19]—incur prohibitive computational costs, limiting their practicality.
To address these challenges, we propose MindGYM, a new framework illustrated in Figure 1 that enhances VLMs via synthetic self-challenging questions with three core stages: (1). Seed Single-Hop Question Synthesis: Generates cognitive questions across text-only (e.g., logical deduction) and multimodal contexts (e.g., scientific diagrams), spanning eight semantic topics like causal reasoning and ethical analysis. (2). Challenging Multi-Hop Question Synthesis: Combines seed questions using cognitive operators—textual strategies (Bridging, Comparison, Temporal) and multimodal reasoning types (Visual-Textual Alignment, Spatial Reasoning, etc.)—to create layered, multi-step problems demanding deeper cognitive engagement. (3). Thinking-Induced Curriculum Fine-Tuning: A structured training pipeline progresses from scaffolded reasoning (e.g., answer prediction with explicit thought paths) to standalone inference, embedding advanced reasoning capacities into model parameters. All these processes are operated on the model to be tuned, via cognitive data self-synthesis and self-challenging training.
Extensive empirical evaluations across 7 benchmarks demonstrate that MindGYM consistently outperforms prior methods on both text-based and multimodal scenarios, achieving +16% gains on challenging tasks like MathVision-Mini with merely 400 synthesis samples, while reducing computational costs in both training and inference time. Notably, MindGYM gains robust generalization across different synthesis sources, and enhanced reasoning depth and breadth via GPT-based scoring analysis over state-of-the-art (SOTA) competitors (+15.77% win rates), enabling the model to handle open-ended tasks requiring multi-modal, high-order cognition.
Our main contributions can be summarized as follows:
  • We propose a new cognitive-grounded synthetic data synthesis pipeline, that systematically generates complex questions considering thinking depth, thinking width, and cross-modality.
  • We show the feasibility of self-challenging evolution for VLMs, with high data-efficiency and progressive training that absorbs explicit human-like reasoning patterns.
  • We verify MindGYM’s effectiveness, practical efficiency, and contextual adaptability with extensive experiments. Our code and data are released at https://github.com/modelscope/data-juicer/tree/MindGYM to foster future progress in vision-language reasoning.
Figure 1: The proposed MindGYM framework architecture consists of three core stages: ① Seed Single-Hop Question Synthesis, ② Challenging Multi-Hop Question Synthesis, and ③ Thinking-Induced Curriculum Fine-Tuning. The process begins with cognitive-aware self-synthesis across text (e.g., mathematical/biological questions) and multimodal (text+image) contexts. Seed questions are then combined via text-based strategies (Bridging, Comparison, Temporal) and multimodal reasoning types (Visual-Textual Alignment, Spatial Reasoning, Causal Inference, Contextual Synthesis) to form self-challenging questions. Finally, a curriculum-driven fine-tuning phase injects structured cognitive competence into the model through masked-task learning. Color-coded regions and arrows illustrate the hierarchical progression from simple reasoning to advanced problem-solving, emphasizing conceptual depth and modal integration.
@@ 2 Related Works
Large vision-language models (VLMs) critically depend on high-quality instruction-tuning datasets to achieve robust performance [imgdiff, 7]. Early efforts, such as OK-VQA [25] and ScienceQA [22], rely on human-annotated data to ensure accuracy and relevance. However, these manual approaches are resource-intensive and difficult to scale. To address this limitation, recent works have explored self-instruct paradigms that automate dataset creation. While text-centric methods like [39, 34, 42] emphasize diversity and complexity in textual instructions, their extension to multimodal domains remains challenging. Multimodal self-instruction frameworks, such as MMInstruct [21] and MMEvol [24], attempt to bridge this gap by iteratively generating image-text pairs. Nevertheless, these methods often suffer from modality inconsistency, where generated textual instructions lack tight visual grounding or rely on shallow cross-modal alignments (e.g., generic image descriptions). This issue arises due to an over-reliance on LLMs for text generation, which lack explicit mechanisms for visual reasoning. Consequently, the resulting datasets fail to challenge VLMs on fine-grained vision-language understanding tasks.
Reasoning-Aware Question Synthesis.
Synthetic question generation has shown success in domains like mathematical reasoning [20, 36], where structured problem spaces enable systematic scaling of complexity [32] and skill decomposition [3]. However, extending these techniques to multimodal domains poses unique challenges due to three key factors: (1) Heterogeneous Semantics: Vision-language reasoning requires handling unstructured inputs with diverse modalities, unlike the formal grammars available in mathematical problems. (2) Cross-Modal Synergy: Current methods focus on single-modality skill progression rather than fostering interactions between modalities. (3) Open-Ended Challenges: Template-driven approaches, such as Visual Program Induction [12], struggle to generate open-ended reasoning tasks that demand higher-order cognitive skills. As a result, existing methods exhibit limited generalizability to complex vision-language tasks, underscoring the need for more flexible and domain-agnostic synthesis frameworks.
Reasoning-Aware Self-Improvement.
Enhancing reasoning capabilities in VLMs often incurs significant computational overhead. RL-based approaches [1] require extensive policy optimization over large-scale reasoning trajectories [14], making them computationally expensive. Iterative refinement techniques [19] further exacerbate costs by introducing multi-stage processing pipelines. Even parameter-efficient methods, such as Chain-of-Thought prompting [40], suffer from latency penalties due to lengthy sequential decoding. Recent advances in test-time scaling [26] demonstrate promising performance gains but at the expense of increased inference latency. These limitations highlight the need for more efficient strategies to improve reasoning without compromising computational feasibility.
Our Position.
MindGYM addresses these three critical limitations of prior art: (1) Modality-inconsistency. With multiple cognitive relationship enhancement in multi-hop questions, we enforce tight semantic coupling in vision-language concepts. (2) Generalization. By adopting domain-agnostic prompts, we explicitly incorporate eight vision meta-topics beyond math, and adaptively increase visual grounding and linguistic complexity on the synthesis data. (3) Computational efficiency. By injecting complex, self-challenging questions directly into the training data, we encourage models through curriculum fine-tuning to perform deep and wide reasoning within a single inference step, eliminating the need for iterative adjustments.
@@ 3 The Proposed MindGYM
In this work, we aim to embed advanced reasoning capabilities into VLMs through synthesized fine-tuning datasets. For large models, conventional single-hop question-answer pairs are insufficient to stimulate reasoning in both depth and breadth. To address this, we propose a progressive self-synthesis framework, outlined in Figure 1, operating through three interconnected stages:
(1) Seed Question Synthesis: Starting with a base VLM $\pi$ (e.g., Qwen-2.5-VL [4]), the model self-generates logically related single-hop questions $\{q^{1}_{i},\ldots,q^{k}_{i}\}_{i\in N}$ to explore diverse reasoning methodologies and topics. Here, $k$ denotes the number of seed questions per context and $N$ represents the number of synthesized data.
(2) Self-Challenging Question Synthesis: Using $\{q^{1}_{i},\ldots,q^{k}_{i}\}$ as inputs, the model constructs structured complex questions $\{(Q_{i},A_{i},T_{i})\}_{i\in N}$ via multi-hop combinations. Each $Q_{i}$ is formed as $Q_{i}=q^{1}_{i}\oplus\ldots\oplus q^{k}_{i}$, where $\oplus$ represents cognitive operators (e.g., composition, contradiction, abstraction) that create hierarchical reasoning structures. The model concurrently generates explicit thinking trajectories $T_{i}$, which documents the rationale behind $Q_{i}$ and $A_{i}$.
(3) Self-Challenging Cognitive Training: The model is fine-tuned on four structured datasets $\{D^{j}_{\text{train}}\}_{j=1}^{4}$ through a curriculum pipeline. Training focuses on predicting masked components across questions $(Q_{i})$, answers $(A_{i})$, and their associated thought paths $(T_{i})$, thereby cultivating deeper and broader reasoning capacity. This adversarial questioning approach systematically improves the model’s ability to navigate ambiguous, multi-step reasoning tasks.
@@
The proposed framework enhances generalization by exposing models to diverse cognitive challenges while injecting interpretable thinking patterns that mirror human-like problem-solving dynamics into the model parameters. Detailed design and implementations are provided in subsequent subsections.
@@ 3.1 Seed Single-Hop Question Synthesis
3.1.1 Diverse Thinking Methodologies
Cognitive-Aware Synthesis Templates.
We adopt a synthesis template that injects thinking elements, consisting of three parts: cognitive setting, background conditions, and synthesis goals: $P=\{P_{cog},P_{context},P_{task}\}$. $P_{context}$ and $P_{task}$ have different instantiations in different synthesis scenarios (single-hop, multi-hop QA, text, multimodal) in our method  [^1] The full prompt is presented in the Appendix B., while $P_{cog}$ remains unchanged to encourage the model to consistently produce responses that are strongly related to thinking.
We first describe the single-hop question scenario, with other tasks detailed later. Based on the given $P_{context}$, $P_{task}$ requires $\pi$ to generate $k$ logically connected seed single-hop questions $q^{1}_{i},...q^{k}_{i}$, at the same time, the model will be explicitly noted that these seed questions will be used later as more complex multi-hop combinations (the specific combination types will be detailed in later Sec.3.2).
Adaptive Stream of Consciousness.
We draw inspiration from a thinking protocol named Thinking-Claude [44] and adapt it to shape $P_{cog}$ as a collection of many diverse thinking processes, (e.g., making and challenging hypotheses, exploring and balancing different perspectives). This makes the model mimic the human thinking process in a layered, context-aware style, aiming for a broader coverage in thinking methodologies. Given different $P_{context}$ and $P_{task}$, $\pi$ dynamically adjusts thinking dimensions, such as problem depth, interpretation angles, and validation rigor, and explicitly expose models’ “minds” dealing with complex scenarios. This adaptive exploration not only encourages the model to incorporate diverse reasoning paths for generalizing to reasoning tasks, but also injects structured variability into training data, countering the memorization biases in conventional pipelines. Empirical validation of $P_{cog}$ is presented in Section 4.3.
3.1.2 Diverse Thinking Objects
Meta-Topics.
We consider both text-only and text-image QA synthesis. Here we first introduce the $P_{context}$ for the text-only case. Building on the cognitive architecture theory [28] and the multi-compartmental knowledge representation [33], we establish eight orthogonal meta-topics spanning four foundational reasoning dimensions: quantitative, causal, temporal, and social-ethical. Unlike prior works focused on limited domain combinations [2], our approach emphasizes broad cognitive coverage:
  • Historical Events: Analyzing past events and their consequences.
  • Economic Principles: Studying resource allocation and valuation.
  • Psychological Insights: Exploring human behavior and cognition.
These eight topics are selected for two core reasons: (1) alignment with widely studied reasoning domains (e.g., mathematics, logic) [17], and (2) extension to synergize with VLMs’ capabilities and potential [30], balancing foundational reasoning with subsequent self-challenges.
By polling one of these meta-topics and their descriptions, we require $\pi$ to self-generate background document as $P_{context}$, then self-generate $k=5$ seed single-hop questions per inference. Empirical results in Section 4.3 demonstrate that this topic-balanced sampling strategy enhances thinking breadth and significantly improves end-to-end model performance compared to unbalanced alternatives. This process allows $\pi$ to explore maximal topic diversity, leveraging its internal knowledge to identify domains and skills requiring advanced reasoning.
Anchoring Image Sources.
@@ 3.1.2 Diverse Thinking Objects
To enhance diversity in seed question generation, MindGYM further applies semantic vectorization to textual components of data in the synthesis pool. We use an empirically determined cosine similarity threshold to evaluate the overlap between newly generated and existing samples. If similarity exceeds this threshold, the system initiates a regeneration cycle to eliminate semantically redundant content, continuing until the seed question pool reaches the target size $N$. This mechanism widens the breadth of reasoning in synthetic data, better preparing it for diverse downstream scenarios.
Unlike prior methods concentrating on domain-specific reasoning (e.g., mathematical tasks), MindGYM expands reasoning scope across three dimensions: cognitive-aware thinking methodologies, image-text interactions, and semantic redundancy filtering. In Section 4.2, we empirically validate MindGYM ’s robustness and effectiveness across varied synthesis sources—including model’s internal knowledge, multilingual contexts, and anchored visual content.

### v2

# MindGYM: What Matters in Question Synthesis for Thinking-Centric Fine-Tuning?

@@ Abstract
Large foundation models face challenges in acquiring transferable, structured thinking abilities, especially when supervised with rigid templates or crowd-annotated instruction datasets. Unlike prior approaches, we focus on a thinking-centric data synthesis paradigm that enables models to evolve through self-generated, cognitively guided data. We propose MindGYM, a structured and scalable framework for question synthesis, composed of: (1) Cognitive Thinking Process Injection, which infuses high-level reasoning objectives to shape the model’s synthesis behavior; (2) Seed Single-Hop Question Synthesis, generating atomic questions from diverse semantic types to encourage broader thinking; and (3) Challenging Multi-Hop QA Synthesis, composing more complex multi-hop questions based on QA seeds for deeper reasoning. Detailed analysis shows that synthetic data generated by our method achieves 16.7% higher average quality and 67.91% lower quality variance compared to baseline sources, highlighting that both high-quality and self-contained data are essential for effective, thinking-oriented finetuning. MindGYM improves performance on six reasoning benchmarks, achieving gains of up to 16% on MathVision using only 400 data samples, and generalizable improvements across different model sizes and architectures. MindGYM underscores the viability of self-challenging mechanisms in refining large model capabilities while minimizing human intervention and resource demands. Code and data are released to promote data-centric research into self-evolving foundation models driven by their internal reasoning capabilities.
11footnotetext: Equal contribution.
@@ 1 Introduction
Large foundation models have emerged as key assistants for tasks requiring complex understanding across tasks and modalities qwen25vl ; qin2024synergydatamultimodallarge . However, enabling robust performance with transferable and efficient thinking abilities remains challenging dj2 ; djsandbox . Manually curated instruction datasets like OK-VQA okvqa and ScienceQA lu2022learn are labor-intensive to scale, while self-supervised synthetic methods such as MMInstruct liu2024mminstruct and MMEvol luo2024mmevol suffer from limited generalization across modality and task types, often failing to produce logically consistent or cognitively diverse data. Meanwhile, reasoning enhancement methods—such as reinforcement learning (RL) akyurek2023rl4f or iterative prompting kumar2024training —incur prohibitive computational costs, limiting their practicality.
To tackle these limitations, we introduce MindGYM, a thinking-centric data synthesis framework aimed at enhancing the cognitive capacity of large models. Rather than relying on task-specific templates or crowd-sourced samples, our approach embeds structured thinking traits into the synthesis process — enabling models to self-generate data that target their cognitive bottlenecks. The proposed framework consists of three key components illustrating in Figure 1:
  1. 1.
    Cognitive Thinking Injection: We infuse structured thinking priors — such as breadth (cross-topic linkage), depth (multi-step deduction), and progression (difficulty scaling) — into the prompt design, guiding generation toward cognitively rich and pedagogically effective samples.
  2. 2.
    Seed Single-Hop Question Synthesis: We synthesize a set of semantically grounded, single-hop QA primitives that span diverse categories including arithmetic, logic, ethics, and causality. These act as composable base units for constructing multi-hop cognitive challenges.
  3. 3.
    Challenging Multi-Hop QA Synthesis: By composing seed questions using thinking-centric operators such as Bridging, Comparison, and Temporal, we generate challenging multi-hop questions that demand higher-order inference and cross-domain understanding.
We validate MindGYM on six reasoning-intensive benchmarks across multiple vision-language models (VLMs) architectures. As shown in Table 1, with only 400 synthetic samples, our method yields significant gains: for example, our method boosts performance of Qwen2.5-VL-7B qwen25vl by over 16% on MathVision-Mini. These improvements are consistent across models of varying scales and architectures, demonstrating the generality of our approach. Further analysis via Data-Juicer dj2 reveals that on Qwen2.5-VL-32B qwen25vl , our synthetic data achieves a +16.7% average quality improvement over baselines and a 67.91% reduction in quality variance. This highlights an important insight: not only does cognitively guided synthesis yield higher-quality data, but stability in quality—as measured by low variance—is critical for consistent fine-tuning. Moreover, beyond the high-quality textual data synthesis presented in the main page, we also explore MindGYM’s scalability to multimodal settings in Appendix C, demonstrating its effectiveness and potential to both VLMs and large language models (LLMs).
To summarize, our contributions are three-fold:
  • We propose MindGYM, a scalable, model-dependent self-synthesis framework for thinking-centric data, which injects structured cognitive priors into both single-hop and multi-hop question generation, encouraging both broad and deep reasoning.
  • We present an in-depth data analysis, showing that reducing variance is as critical as improving mean quality for stable fine-tuning. The proposed method generates cognitively guided data that excels on both fronts.
  • We demonstrate that self-challenging data evolution significantly enhances reasoning performance across six benchmarks, even under limited supervision, and generalizes well across different models. All code and data are released at https://github.com/modelscope/data-juicer/tree/MindGYM/ to shed light on further research and applications.
@@ 2 Related Works
High-quality instruction data is pivotal for vision-language models (VLMs) to align with human intent jiao2024imgdiffcontrastivedatasynthesis ; kddtutorial . While early benchmarks like OK-VQA okvqa and ScienceQA lu2022learn rely on costly human annotation, recent work embraces self-instruction paradigms to scale data generation wang2023self ; taori2023alpaca ; xu2023wizardlm . In the multimodal domain, MMInstruct liu2024mminstruct and MMEvol luo2024mmevol extend this idea, but often suffer from modality inconsistency, where visual inputs are weakly grounded to textual outputs. These limitations stem from over-reliance on LLMs without structured visual reasoning. Our work departs from these trends by embedding cognitively structured priors into prompt design, generating semantically aligned, cross-modal instruction data.
Thinking-Centric Question Synthesis
Synthetic data generation has long been a tool for improving model reasoning. In language-only domains, question synthesis approaches have emphasized structure, difficulty scaling, and multi-step deduction, particularly in math or logic domains lewkowycz2022solving ; shao2024deepseekmath . However, most existing methods either depend on rigid templates or domain-specific structures, which constrains their generalizability. Multimodal question synthesis introduces further complexity due to the need for cross-modal reasoning and abstract concept grounding. Methods like Visual Program Induction pmlr-v162-duan22c or GRILL parcalabescu2022valse attempt structured generation but often produce shallow or repetitive questions. Moreover, many pipelines prioritize syntactic fluency over cognitive depth, failing to elicit the higher-order thinking needed for robust generalization. This motivates the need for synthesis frameworks that directly embed cognitive traits—such as reasoning breadth, depth, and progression—into the data itself.
Self-Improving Reasoning with Limited Supervision
Self-improvement paradigms for reasoning have emerged to reduce human intervention and labeling costs. Reinforcement learning–based methods (e.g., RL4F akyurek2023rl4f ) and iterative self-refinement kumar2024training fine-tune models through repeated exploration and feedback loops but require significant computational resources. Lightweight techniques like Chain-of-Thought prompting wei2022chain and Tree-of-Thoughts yao2023tree improve test-time reasoning but often rely on long sequences or search-time strategies, which are inefficient for deployment. In contrast, data-centric self-improvement strategies aim to modify the training signal itself rather than the inference process. MindGYM fits into this paradigm by synthesizing cognitively aligned, reasoning-targeted samples that directly support robust and efficient learning, without the need for test-time decoding tricks or expensive policy optimization.
@@ 3 The Proposed MindGYM
Figure 1: The proposed MindGYM framework incorporates a cognitively guided data synthesis pipeline with four stages: ① Context Generation, ② Single-Hop Question Synthesis, ③ Multi-Hop Composition with Thinking Trace, and ④ Structured Extraction. Starting from a meta-topic and guided by a shared cognitive objective, the model iteratively builds background context, atomic reasoning steps, composite questions with interpretable thinking traces, and final structured QA samples for downstream use. Color-coded regions and arrows illustrate the hierarchical progression from simple reasoning to advanced problem-solving and emphasizing conceptual depth.
@@
In this work, we aim to embed advanced thinking capabilities into large models through self-synthesized fine-tuning datasets. Prior research has shown that conventional single-hop question–answer pairs are often inadequate to stimulate thinking along both depth and breadth dimensions, especially for tasks that require multi-step deduction or complex reasoning. To address this, we introduce a progressive self-synthesis framework, outlined in Figure 1, operating through three interconnected components: (1) Cognitive Thinking Injection, where diverse human-like reasoning strategies are embedded into the synthesis process via structured prompts; (2) Seed Single-Hop Question Synthesis, where the model generates logically grounded, cognitively-aware atomic questions; and (3) Challenging Multi-Hop QA Synthesis, where these atomic units are composed into complex, multi-step questions with explicit reasoning trajectories.
@@
This framework enhances generalization by exposing models to diverse cognitive challenges while injecting interpretable thinking patterns that mirror human-like problem-solving dynamics into the model parameters. Detailed designs and implementations are provided in subsequent subsections.
3.1 Cognitive Thinking Injection
Cognitive-Aware Synthesis Templates.
[…]
We propose a four-stage cognitive thinking injection framework that incrementally builds structured reasoning traces using a unified generation model $\pi$, guided by $P_{\text{cog}}$ and stage-specific $P_{\text{task}}$, with Stage 1 and Stage 2 described in Section 3.2, and Stage 3 and Stage 4 in Section 3.3:
  • Stage 1: Background Context Generation. Given a meta-topic, the model generates a background passage $P_{\text{context}}$ using $P_{\text{cog}}$ and a context prompt $P_{\text{task}}^{\text{context}}$: $\pi(P_{\text{cog}},P_{\text{task}}^{\text{context}})\rightarrow P_{\text{context}}.$
  • Stage 2: Single-Hop Question Generation. Conditioned on $P_{\text{cog}}$, $P_{\text{context}}$, and $P_{\text{task}}^{\text{single-hop}}$, the model generates $k$ logically independent seed questions: $\pi(P_{\text{cog}},P_{\text{context}},P_{\text{task}}^{\text{single-hop}})\rightarrow\{q_{1},\dots,q_{k}\}.$
  • Stage 3: Cognitive Composition with Adaptive Types. Using the single-hop questions and a multi-hop prompt, the model composes a complex multi-hop question and a corresponding reasoning trace: $\pi(P_{\text{cog}},P_{\text{context}}^{\text{single-hop}},P_{\text{task}}^{\text{multi-hop}})\rightarrow(q_{\text{multi}},\texttt{Thinking}).$
  • Stage 4: Structured Extraction. Given the multi-hop results and a structured prompt, $\pi$ produces the final schema-aligned data: $\pi(P_{\text{cog}},P_{\text{context}}^{\text{multi-hop}},P_{\text{task}}^{\text{structured}})\rightarrow(\texttt{Question},\texttt{Answer},\texttt{Others}).$
Adaptive Stream of Consciousness.
To implement $P_{\text{cog}}$, we adapt the Thinking-Claude protocol claude , structuring it as a flexible suite of cognitive operations (e.g., hypothesis formulation and testing, multi-angle interpretation, counterfactual reasoning, and self-verification). These thinking elements are modular and reusable, enabling the model to emulate layered, human-like thinking. Across stages, $\pi$ is prompted to verbalize and externalize its internal thinking steps, thereby producing interpretable structured traces that reveal how the model navigates complex problem spaces.
This cognitively guided synthesis approach enhances generalization by exposing the model to diverse reasoning challenges while injecting structured and interpretable thinking into training data. The resulting dataset contains rich cognitive signals that are later leveraged during fine-tuning (see Section 3.4). Empirical evidence for the efficacy of $P_{\text{cog}}$ is presented in Section 4.4.
@@ 3.2 Seed Single-Hop Question Synthesis
Cognitively Grounded Meta-Topics.
To ensure reasoning diversity and depth, we construct eight orthogonal meta-topics derived from established cognitive theories, including unified cognitive architectures newell1994unified and dual-representation models sun2001duality . These meta-topics span four core reasoning dimensions: quantitative, causal, temporal, and social-ethical, offering a comprehensive blueprint for cognitively rich question generation. Detailed meta-topics are provided in Appendix B.2.
Stage 1: Background Context Generation.
Given a sampled meta-topic and its brief description, the model is prompted with a cognitive grounding prompt $P_{\text{cog}}$ and a context-specific instruction $P_{\text{task}}^{\text{context}}$ to self-generate a short but semantically rich background passage $P_{\text{context}}$. This passage serves as a reference for subsequent single-hop question generation.
Stage 2: Single-Hop Question Generation.
Conditioned on the generated background passage $P_{\text{context}}$, the cognitive intent $P_{\text{cog}}$, and a task-specific instruction $P_{\text{task}}^{\text{single-hop}}$, the model generates a small batch (up to $k=5$) of logically grounded, atomic single-hop questions $\{q_{1},q_{2},\dots,q_{k}\}$. These questions are designed to probe individual reasoning steps based on the passage and are explicitly treated as intermediate components for later multi-hop composition in Section 3.3.
@@
To enhance diversity in seed question generation, MindGYM further applies semantic vectorization to textual components of data in the synthesis pool. We use an empirically determined cosine similarity threshold to evaluate the overlap between newly generated and existing samples. If similarity exceeds the threshold, the system initiates a regeneration cycle to eliminate semantically redundant content, continuing until the seed question pool reaches the target size $N$. This reject sampling mechanism widens the breadth of reasoning in synthetic data.
Unlike prior works that narrowly focus on limited task domains austin2021program ; huang2023towards (e.g., arithmetic, logic puzzles), MindGYM emphasizes (1) internalized background grounding, (2) cognitively structured topic guidance, and (3) semantic diversity enforcement. In Section 4.4, we empirically validate MindGYM robustness and effectiveness across varied synthesis sources, including model’s internal knowledge and multimodal contexts.